OpenStudy (anonymous):

Explain how you would prove if four given points are coplanar. Use your method to determine if A (3, 4, -2), B (8, 5, 0), C (1, 10, -6) and D (9, 2, 2) are coplanar.

OpenStudy (unklerhaukus):

with a matrix?

OpenStudy (unklerhaukus):

for 4 points to be co-planar , some of them must be linear combinations of others

OpenStudy (anonymous):

and how should i proceed please.

ganeshie8 (ganeshie8):

or you could work it by finding the distance between one of the points to the plane determined by remaining 3 points

ganeshie8 (ganeshie8):

are u familiar with cross product ?

OpenStudy (anonymous):

im not sure, im doing the xproduct right

OpenStudy (anonymous):

AB=(5,1,2) AC=(-2,6,-4) AD=(6,-2,-0) for ABxAC ive got (5,-1,10) Im stuck here dont know how to go foward

ganeshie8 (ganeshie8):

check your AD again

OpenStudy (anonymous):

yeah its (-6,-2,4)

ganeshie8 (ganeshie8):

A (3, 4, -2), B (8, 5, 0), C (1, 10, -6) and D (9, 2, 2) AB=(5,1,2) AC=(-2,6,-4) AD=(6,-2, 4) work ABxAC again

ganeshie8 (ganeshie8):

\[\large AB\times AC = \left|\begin{array}{|c|c|c|} i&j&k\\5&1&2\\-2&6&-4\end{array}\right| = -16i + 16j +32k\]

ganeshie8 (ganeshie8):

so that the normal vector the plane determined by lines AB and AC

ganeshie8 (ganeshie8):

ABxAC = (-16, 16, 32) dot this with AD, if both are perpendicular then the dot product will be 0 confirming that D lies in the same plane as A,B,C

ganeshie8 (ganeshie8):

AD . (AB x AC) = (6,-2, 4) . (-16, 16, 32) = ?

OpenStudy (anonymous):

yeah its zero

OpenStudy (anonymous):

thx for you help, I will go and review the cross product.. forgot everything about that

ganeshie8 (ganeshie8):